Shift report 20000722 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Jul 22 2000 - 00:08:00 EDT

  • Next message: http account: "Shift report 20000722 00:00-08:00"

    ------------------------------------------------------------
    Date: Sat Jul 22 00:07:59 2000, Shift: 16:00-24:00
    Supervisor: YKL, with: TKO
    Log book#3 pages 42 - 43
    
    Shift Summary:
    ------------------------------------------------------------
    Trouble with PASS system at 2 o'clock;Booster Main Magnet 
    trip; mostly working for "survival" of injection in 
    their words. Finally both blue and yellow beam intensity
    at 2000 x E6 for over 40 minutes but no ramping.
    No data were taken.TPC anodes were turned off throughout.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Jul 22 2000 - 00:11:33 EDT